# OpenAD: Open-World Autonomous Driving Benchmark for 3D Object Detection

### v1

@@ Abstract
Open-world autonomous driving encompasses domain generalization and open-vocabulary. Domain generalization refers to the capabilities of autonomous driving systems across different scenarios and sensor parameter configurations. Open vocabulary pertains to the ability to recognize various semantic categories not encountered during training. In this paper, we introduce OpenAD, the first real-world open-world autonomous driving benchmark for 3D object detection. OpenAD is built on a corner case discovery and annotation pipeline integrating with a multimodal large language model (MLLM). The proposed pipeline annotates corner case objects in a unified format for five autonomous driving perception datasets with 2000 scenarios. In addition, we devise evaluation methodologies and evaluate various 2D and 3D open-world and specialized models. Moreover, we propose a vision-centric 3D open-world object detection baseline and further introduce an ensemble method by fusing general and specialized models to address the issue of lower precision in existing open-world methods for the OpenAD benchmark. Data, toolkit codes, and evaluation codes are released at https://github.com/VDIGPKU/OpenAD.
@@ 1 Introduction
Table 1: Open-world autonomous driving datasets or benchmarks. “*” means rough estimates. OpenAD is the first real-world open-world benchmark for autonomous driving 3D perception. Compared to other real-world datasets, OpenAD boasts a richer variety of categories and more instances.
@@
Synthetic Fire Hydrants [8] Cam. 30,000 1 30.0K* Bbox(2D)
Synthetic Crosswalks [8] Cam. 20,000 1 20.0K* Bbox(2D)
CARLA-WildLife [45] Cam. Depth 26 18 65* Inst. mask(2D)
MUAD [20] Cam. Depth 4,641 9 30.0K Sem. mask(2D)
@@
With the rapid development of autonomous driving systems, open-world perception has garnered significant and growing attention from the research community. Open-world perception endeavors to develop a model that exhibits robust performance across novel domains, diverse sensor configurations, and various corner case objects. The two most pivotal factors in open-world perception are domain generalization and open-vocabulary.
Domain generalization refers to the performance of a model when confronted with new scenarios outside the training domain. It is a crucial issue that must be addressed to achieve Level 4 autonomous driving. Within autonomous driving 3D perception, the current methodologies [29, 1] for evaluating scenario generalization entail training on a specific dataset and then transferring the trained model to a distinct dataset for subsequent testing.
Open-vocabulary denotes the recognition capability of perception models toward semantic categories that are not present or unlabeled within the training domain. Open-vocabulary perception serves as the foundation for subsequent inference and planning in autonomous driving systems. For instance, determining whether an object is collidable, whether it might suddenly move, or whether it signifies that certain surrounding areas are not traversable, necessitates an accurate semantic description of the object in the first place.
Many works are proposed to address these two issues. However, researchers meet three challenges when developing open-world perception models. The first challenge in 3D open-world perception for autonomous driving lies in the scarcity of evaluation benchmarks. Specifically, a unified benchmark for domain transfer evaluation is currently absent, and due to the varying formats of individual datasets, researchers must expend considerable effort on the engineering aspect of format alignment. Besides, the current 3D perception datasets possess a limited number of semantic categories, lacking effective evaluation for current open-vocabulary 3D perception models.
The second challenge is the difficulty in training open-world perception models due to the limited scales of publicly available 3D perception datasets. Though some open-world natural language models and 2D perception models have recently leveraged large-scale Internet data for training. How to transfer these models’ capabilities or 2D data to 3D open-world perception is an important and timely research problem.
The last challenge is the relatively low precision of existing open-world perception models. While specialized models trained on autonomous driving perception datasets lack the capability to generalize to the open world, they exhibit stronger predictive power for seen categories and achieve good performance. This indicates that, as the specialized models, the low precision of open-world perception models limits their real-world application. Consequently, current open-world perception models cannot yet replace specialized models in practice.
To address the aforementioned challenges, we propose OpenAD, an Open-World Autonomous Driving Benchmark for 3D Object Detection. We align the format of five existing autonomous driving perception datasets, select 2,000 scenes, annotate thousands of corner case objects with MLLMs, and develop open-world evaluation metrics to overcome the first challenge of scarcity of evaluation benchmarks. Then, we introduce a vision-centric 3D open-world object detection baseline by utilizing existing 2D open-world perception models to resolve the second challenge. Finally, we further design a fusion method to address the last challenge by leveraging the strengths of open-world perception models (or general models) and specialized models to improve the 3D open-world perception results.
@@
  • We propose an open-world benchmark that simultaneously evaluates object detectors’ domain generalization and open-vocabulary capabilities. To our knowledge, this is the first real-world autonomous driving benchmark for 3D open-world object detection.
@@
  • We propose a baseline method for 3D open-world perception by combining 2D open-world models. Besides, we analyze the strengths and weaknesses of open-world and specialized models, and further introduce a fusion approach to leverage both advantages.
@@ 2.1 Benchmark for Open-world Object Detection
2D Benchmark. Various datasets [39, 24, 53, 35, 21] has been used for 2D open-vocabulary object detection evaluation. The most commonly used one is LVIS dataset [24], which contains 1,203 categories.
In the autonomous driving area, as shown in Table 1, many datasets [26, 8, 45, 20, 13, 22, 50, 5, 26, 34] has been proposed too. Among them, CODA [34] is a road corner case dataset for 2D object detection in autonomous driving with 1,500 road driving scenes containing bounding box annotations for 34 categories. However, some datasets only provide semantic segmentation annotations without specific instances or annotate objects as abnormal but lack semantic tags. Moreover, datasets collected from real-world driving data are on a small scale, while synthetic data from simulation platforms such as CARLA [18] lacks realism, making it difficult to conduct effective evaluations. In contrast, our OpenAD offers large-scale 2D and 3D bounding box annotations from real-world data for a more comprehensive open-world object detection evaluation.
3D Benchmark. The 3D open-world benchmarks can be divided into two categories: indoor and outdoor scenarios. For indoor scenarios, SUN-RGBD [54] and ScanNet [17] are two real-world datasets often used for open-world evaluation, containing about 700 and 21 categories, respectively. For outdoor or autonomous driving scenarios, AnoVox [6] is a synthetic dataset containing instance masks of 35 categories for open-world evaluation. However, due to limited simulation assets, the quality and instance diversity of the synthetic data are inferior to real-world data. In addition to AnoVox, existing real-data 3D object detection datasets for autonomous driving [9, 46, 58, 55, 21] only contain a few object categories, which can hardly be used to evaluate open-world models. To address this issue, we propose OpenAD, which is constructed from real-world data and contains 206 different corner-case object categories that appeared in autonomous driving scenarios.
@@
To address the out-of-distribution (OOD) or anomaly detection, earlier approaches [64] typically employed decision boundary, clustering, and so forth, to discover OOD objects. Recently methods  [31, 71, 44, 42, 60, 62, 68, 35, 56, 70, 15, 59, 23] employ text encoders, i.e. CLIP [52], to align text features of corresponding category labels with the box features. Specifically, OVR-CNN [68] aligns the image features with caption embeddings. GLIP [35] unifies object detection and phrase grounding for pre-training. OWL-ViT v2 [47] uses a pretrained detector to generate pseudo labels on image-text pairs to scale up detection data for self-training. YOLO-World [15] adopts a YOLO-type architecture for open-vocabulary detection and achieves good efficiency. However, all these methods require predefined object categories during inference.
More recently, some open-ended methods [16, 66, 40] propose to utilize natural language decoders to provide language descriptions, which enables them to generate category labels from RoI features directly. More specifically, GenerateU [16] introduces a language model to generate class labels directly from regions of interest. DetClipv3 [66] introduced an object captioner to generate class labels during inference and image-level descriptions for training. VL-SAM [40] introduces a training-free framework with the attention map as prompts.
@@
In contrast to 2D open-world object detection tasks, 3D open-world object detection tasks are more challenging due to the limited training datasets and complex 3D environments. To alleviate this issue, most existing 3D open-world models bring power from pretrained 2D open-world models or utilize abundant 2D training datasets.
For instance, some indoor 3D open-world detection methods like OV-3DET [43] and INHA [30] use a pretrained 2D object detector to guide the 3D detector to find novel objects. Similarly, Coda [10] utilizes 3D box geometry priors and 2D semantic open-vocabulary priors to generate pseudo 3D box labels of novel categories. FM-OV3D [69] utilizes stable diffusion to generate data containing OOD objects. As for outdoor methods, FnP [19] uses region VLMs and a Greedy Box Seeker to generate annotations for novel classes during training. OV-Uni3DETR [57] utilizes images from other 2D datasets and 2D bounding boxes or instance masks generated by an open-vocabulary detector.
However, these existing 3D open-vocabulary detection models require predefined object categories during inference. To address this issue, we introduce a vision-centric open-ended 3D object detection method, which can directly generate unlimited category labels during inference.
@@ 3.1 Scenes and Annotation
Figure 2: Data composition of OpenAD. We utilized a greater number of scenes from the ONCE dataset since the other four datasets, sampled from major cities in the United States, share certain similarities, and the Once dataset contains a richer variety of corner case scenarios. Additionally, we annotated each object with an indication of whether its category was observed in the training set of each dataset, allowing for separate evaluations of the model’s specialized performance and open-vocabulary performance.
The 2,000 scenes in OpenAD are carefully selected from five large-scale autonomous driving perception datasets: Argoverse 2 [58], KITTI [21], nuScenes [9], ONCE [46] and Waymo [55], as illustrated in Figure 2. These scenes are collected from different countries and regions, and have different sensor configurations. Each scene has the temporal camera and LiDAR inputs and contains at least one corner case object that the original dataset has not annotated.
For 3D bounding box labels, we annotate 6,597 corner case objects across these 2,000 scenarios, combined with the annotations of 13,164 common objects in the original dataset, resulting in 19,761 objects in total. The location and size of all objects are manually annotated using 3D and 2D bounding boxes, while their semantics categories are labeled with natural language tags, which can be divided into 206 classes. We illustrate some corner case objects in  Figure 1. OpenAD encompasses both abnormal forms of common objects, such as bicycles hanging from the rear of cars, cars with doors open, and motorcycles with rain covers, as well as uncommon objects, including open manholes cover, cement blocks, and tangled wires scattered on the ground.
Concurrently, we have annotated each object with a “seen/unseen” label, indicating whether the categories of the objects have appeared in the training set of each dataset. This label is intended to facilitate the evaluation process by enabling a straightforward separation of objects that the model has encountered (seen) and those it has not (unseen), once the training dataset is specified. Moreover, we offer a toolkit code that consolidates scenes from five original datasets into a unified format, converts them into OpenAD data, and facilitates the loading and visualization process.
@@ 3.2 Evaluation Metrics
OpenAD provides evaluations for both 2D and 3D open-world object detection.
@@
For 3D object detection, the center distance is adopted as the positional score following nuScenes, and we use the same semantic score as the 2D detection task. Similar to nuScenes, we adopt a multi-threshold averaging method for AP calculation. Specifically, we compute AP across 12 thresholds, combining positional thresholds of 0.5m, 1m, 2m, and 4m with semantic similarity thresholds of 0.5, 0.7, and 0.9, and then average these AP values.
@@
In/Out Domain & Seen/Unseen AR. To evaluate the model’s domain generalization ability and open-vocabulary capability separately, we calculate the AR based on whether the scene is within the training domain and whether the object semantics have been seen during training. The positional thresholds for this metric are defined as above, whereas the semantic similarity thresholds are fixed at 0.9.
@@ 4 Construction of OpenAD
OpenAD is inspired by the CODA [34] dataset, which focuses on 2D corner cases in autonomous driving. However, certain objects, such as cables or nails close to the road surface, and signboards hanging on walls, cannot be detected solely by LiDAR. Therefore, unlike CODA’s LiDAR-based pipeline, we propose a vision-centric semi-automated annotation pipeline, as shown in Figure 3.
Figure 3: Annotation pipeline. OpenAD is built on a corner case discovery and annotation pipeline that integrates with a multimodal large language model (MLLM).
We use an MLLM Abnormal Filter to identify scenes containing corner cases within the validation and test sets of five autonomous driving datasets, followed by manual filtering. After that, we annotated the corner case objects with 2D bounding boxes.
For objects with relatively complete 3D geometry formed by point clouds, we adopt a methodology similar to CODA by employing point-cloud clustering algorithms [7]. We then utilize camera parameters to project 2D bounding boxes into the point cloud space and identify the corresponding clusters. Finally, the bounding boxes are manually corrected. For objects that are difficult to detect through point-cloud clustering, we manually annotate 3D bounding boxes by referencing multi-view images.
For category labels, we send images with 2D bounding boxes to an MLLM for semantic annotation and indicate for each object whether its category has been seen in each dataset. To select the best MLLM and prompts for object recognition, we manually select 30 challenging annotated image samples and evaluate the accuracy of each MLLM and prompt. We use GPT-4V [48], Claude 3 Opus [2], and InternVL 1.5 [14], with InternVL exhibiting the best performance. Our experiments also reveal that closed image prompts, such as 2D bounding boxes or circles, yield the best results, whereas marking the object of inquiry on the image with arrows yields slightly inferior results. The final MLLM and prompt achieve an accuracy rate of approximately 65% on the 30 challenging samples and around 90% on the entire data. Objects like open manholes and wires falling on the road are difficult to identify for existing MLLMs.
Note that though we have utilized tools such as MLLM to automate some stages as much as possible to reduce manual workload, we have also incorporated manual verification into each stage to ensure the accuracy of annotations.

### v2

@@ Abstract
Open-world perception aims to develop a model adaptable to novel domains and various sensor configurations and can understand uncommon objects and corner cases. However, current research lacks sufficiently comprehensive open-world 3D perception benchmarks and robust generalizable methodologies. This paper introduces OpenAD, the first real open-world autonomous driving benchmark for 3D object detection. OpenAD is built upon a corner case discovery and annotation pipeline that integrates with a multimodal large language model (MLLM). The proposed pipeline annotates corner case objects in a unified format for five autonomous driving perception datasets with 2000 scenarios. In addition, we devise evaluation methodologies and evaluate various open-world and specialized 2D and 3D models. Moreover, we propose a vision-centric 3D open-world object detection baseline and further introduce an ensemble method by fusing general and specialized models to address the issue of lower precision in existing open-world methods for the OpenAD benchmark. We host an online challenge on EvalAI (2D & 3D). Data, toolkit codes, and evaluation codes are available at https://github.com/VDIGPKU/OpenAD.
@@ 1 Introduction
3D perception, which conforms to practical spatial definitions and physical laws, has become indispensable in autonomous driving systems. In the pursuit of advanced autonomous driving, the necessity for open-world capabilities has been recognized. The two most pivotal factors in open-world perception are domain generalization and being open ended. Domain generalization refers to the performance of a model when faced with new streets, regions or countries, as well as varied vehicle types. Within 3D perception for autonomous driving, existing methodologies [28, 1] for evaluating scenario generalization entail training on a specific dataset and then transferring the trained model to a distinct dataset for subsequent testing. Open-ended denotes the capability to provide specific category descriptions for any common or uncommon instance. Open-ended perception is the foundation for subsequent inference and planning in autonomous driving systems. For instance, determining whether an object is collidable, whether it might suddenly move, or whether it signifies that certain areas are not traversable, necessitates an accurate semantic description of the object in the first place.
Many works have been proposed to address these two issues. However, numerous challenges remain when developing open-world perception models. The first challenge in 3D open-world perception for autonomous driving lies in the scarcity of evaluation benchmarks. Specifically, a unified benchmark for domain transfer evaluation is currently absent, and due to the varying formats of individual datasets, researchers must expend considerable effort on the engineering aspect of format alignment. In addition, current 3D perception datasets possess a limited number of common semantic categories, lacking an effective evaluation for open-ended 3D perception models.
The second challenge is the difficulty of training open-world perception models due to the limited scales of publicly available 3D perception datasets. Some open-world language models and 2D perception models have recently used large-scale Internet data for training. How to transfer these models’ capabilities to 3D open-world perception is an important and timely research problem.
The last challenge is the relatively low precision of the existing open-world perception models. Specialized models, which lack the capability to understand uncommon objects, exhibit stronger performance for common categories. Consequently, current open-world perception models cannot yet replace specialized models in practice.
Table 1: Open-world autonomous driving datasets or benchmarks. “*” means rough estimates. OpenAD is the first real-world open-world benchmark for autonomous driving 3D perception. Compared to other real-world datasets, OpenAD boasts greater category diversity and more instances.
@@
Synthetic Fire Hydrants [7] Cam. 30,000 1 30K* Bbox(2D)
Synthetic Crosswalks [7] Cam. 20,000 1 20K* Bbox(2D)
CARLA-WildLife [44] Cam. Depth 26 18 65 Inst. mask(2D)
MUAD [19] Cam. Depth 4,641 9 30K Sem. mask(2D)
@@
To address the aforementioned challenges, we propose OpenAD, an Open-World Autonomous Driving Benchmark for 3D Object Detection. We align the format of five existing autonomous driving perception datasets, select 2,000 scenes, annotate thousands of corner case objects with MLLMs, and develop open-world evaluation metrics to overcome the scarcity of evaluation benchmarks. Then, we introduce a novel vision-centric framework for 3D open-world perception, which utilizes existing 2D open-world perception models to resolve the second challenge. Compared to existing methods, this approach achieves higher average Precision and Recall on the OpenAD benchmark. Finally, we further design a fusion method to address the last challenge by leveraging the strengths of open-world perception models and specialized models to improve the 3D open-world perception results.
@@
  • We propose an open-world benchmark that simultaneously evaluates object detectors’ domain generalization and open-ended capabilities. To our knowledge, this is the first real-world autonomous driving benchmark for 3D open-world object detection.
@@
  • We propose a novel vision-centric framework for 3D open-world perception. In addition, we analyze the strengths and weaknesses of open-world and specialized models and further introduce a fusion approach to utilize both advantages.
@@ 2.1 Benchmark for Open-world Object Detection
2D Benchmark. Various datasets [38, 23, 51, 34, 20] have been used for 2D open-vocabulary and open-ended object detection evaluation. The most widely used is the LVIS dataset [23], which contains 1,203 categories. In the autonomous driving area, as shown in Table 1, many datasets [25, 7, 44, 19, 12, 21, 48, 4, 25, 33] have been proposed. However, some datasets only provide semantic segmentation annotations without specific instances or annotate objects as abnormal but lack semantic tags. Moreover, datasets collected from real-world driving data are on a small scale, while synthetic data from simulation platforms such as CARLA [17] lack realism, making it difficult to conduct effective evaluations. In contrast, our OpenAD offers large-scale 2D and 3D bounding box annotations from real-world data for a more comprehensive open-world object detection evaluation.
3D Benchmark. The 3D open-world benchmarks can be divided into two categories: indoor and outdoor scenarios. For indoor scenarios, SUN-RGBD [52] and ScanNet [16] are two real-world datasets often used for open-world evaluation, containing about 700 and 21 categories, respectively. For outdoor or autonomous driving scenarios, AnoVox [5] is a synthetic dataset, containing instance masks of 35 categories for open-world evaluation. However, due to limited simulation assets, the quality and instance diversity of the synthetic data are inferior to real-world data. Existing real-data 3D object detection datasets for autonomous driving [8, 45, 56, 53, 20] only contain a few common object categories, which can hardly be used to evaluate open-world models. To address these issues, OpenAD is constructed from real-world data and contains 206 different corner-case object categories that appeared in autonomous driving scenarios.
Additionally, the metrics of existing benchmarks are not designed for open-world detection, as their ground-truth annotations presuppose fixed categories. Evaluation inaccuracies may arise when semantic labels overlap or when open-ended models predict synonymous terms. Furthermore, long-tail objects can disproportionately skew the computation of these metrics. OpenAD has redesigned metrics to address this issue.
@@
To address out-of-distribution (OOD) or anomaly detection, earlier approaches [62] typically employed decision boundaries, clustering, etc., to discover OOD objects. Recent methods  [30, 70, 43, 41, 58, 60, 67, 34, 54, 69, 14, 57, 22] employ text encoders, such as CLIP [50], to align the text features of the corresponding category labels with the box features. Specifically, GLIP [34] unifies object detection and phrase grounding for pre-training. OWL-ViT v2 [46] uses a pretrained detector to generate pseudo labels on image-text pairs to scale up detection data for self-training. YOLO-World [14] adopts a open-vocabulary YOLO-type architecture and achieves good efficiency. However, all these methods require predefined object categories during inference.
More recently, some open-ended methods [15, 65, 39] utilize natural language decoders to provide language descriptions and facilitate generating category labels from RoI features directly. More specifically, GenerateU [15] introduces a language model to generate class labels directly from regions of interest. DetClipv3 [65] introduced an object captioner to generate class labels during inference and image-level descriptions for training. VL-SAM [39] introduces a training-free framework with the attention map as prompts.
@@
In contrast to 2D open-world object detection tasks, 3D open-world object detection tasks are more challenging due to the limited training datasets and complex 3D environments. To alleviate this issue, most existing 3D open-world models leverage power of pretrained 2D open-world models or utilize abundant 2D training datasets.
For instance, some indoor 3D open-world detection methods like OV-3DET [42], INHA [29], and ImOV3D [64] use a pretrained 2D object detector to guide the 3D detector to find novel objects. Similarly, Coda [9] utilizes 3D box geometry priors and 2D semantic open-vocabulary priors to generate pseudo 3D box labels of novel categories. FM-OV3D [68] utilizes stable diffusion to generate data containing OOD objects. As for outdoor methods, FnP [18] uses region VLMs and a Greedy Box Seeker to generate annotations for novel classes during training. OV-Uni3DETR [55] utilizes images from other 2D datasets and 2D bounding boxes or instance masks generated by an open-vocabulary detector.
However, these existing 3D open-vocabulary detection models require predefined object categories during inference. To address this issue, we introduce a vision-centric open-ended 3D object detection method, which can directly generate effectively unlimited category labels during inference.
@@ 3.1 Scenes and Annotation
The 2,000 scenes in OpenAD are carefully selected from five large-scale autonomous driving perception datasets: Argoverse 2 [56], KITTI [20], nuScenes [8], ONCE [45] and Waymo [53], as illustrated in Figure 2. These scenes are collected from different countries and regions, and have different sensor configurations. Each scene has temporal camera and LiDAR inputs and contains at least one corner case object that the original dataset has not annotated.
For 3D bounding box labels, we annotate 6,597 corner case objects across these 2,000 scenarios, combined with the annotations of 13,164 common objects in the original dataset, resulting in 19,761 objects in total. The location and size of all objects are manually annotated using 3D and 2D bounding boxes, while their semantic categories are labeled with natural language tags, which can be divided into 206 classes. We illustrate some corner case objects in Figure 1. OpenAD encompasses both abnormal forms of common objects, such as bicycles hanging from the rear of cars, cars with doors open, and motorcycles with rain covers, as well as uncommon objects, including open manhole covers, cement blocks, and tangled wires scattered on the ground.
In addition, we have annotated each object with a “seen/unseen” label, indicating whether the categories of the objects have appeared in the training set of each dataset. This label is intended to facilitate the evaluation process by enabling a straightforward separation of objects that the model has encountered (seen) and those it has not (unseen), once the training dataset is specified. In addition, we offer a toolkit code that consolidates scenes from five original datasets into a unified format, converts them to OpenAD data, and facilitates the loading and visualization process.
Figure 2: Data composition of OpenAD. OpenAD covers multiple cities in various countries, including scenes during the day and night, on different weather and road scenarios. Additionally, we annotate each object with an indication of whether its category is observed in the training set of each dataset, allowing for separate evaluations of the model’s specialized performance and open-ended performance.
@@ 3.2 Evaluation Metrics
In contrast to existing benchmarks, we utilize natural language annotations and conduct multi-threshold matching against string-formatted model predictions. This evaluation framework not only resolves synonym challenges but also demonstrates stronger consistency with the conceptual framework of Open-World paradigms.
@@
For 3D object detection, the center distance is adopted as the positional score following nuScenes, and we use the same semantic score as the 2D detection task. Similar to nuScenes, we adopt a multi-threshold averaging method for AP calculation. Specifically, it computes AP across 12 thresholds, combining positional thresholds of 0.5m, 1m, 2m, and 4m with semantic similarity thresholds of 0.5, 0.7, and 0.9, and then average these AP values.
@@
In/Out Domain & Seen/Unseen AR. To evaluate the model’s domain generalization ability and open-ended capability separately, we calculate the AR based on whether the scene is within the training domain and whether the object semantics have been seen during training. The positional thresholds for this metric are defined as above, whereas the semantic similarity thresholds are fixed at 0.9.
@@ 4 Construction of OpenAD
We propose a vision-centric semi-automated annotation pipeline for OpenAD, as shown in Figure 3. This differs from the existing LiDAR-based pipeline [33] because certain objects, such as cables or nails close to the road surface and wall-hung signboards, cannot be detected solely by LiDAR.
Figure 3: Annotation pipeline. OpenAD is built upon a corner case discovery and annotation pipeline that integrates with a multimodal large language model (MLLM).
We use an MLLM Abnormal Filter to identify scenes containing corner cases within the validation and test sets of five autonomous driving datasets, followed by manual filtering. We then annotate the corner case objects with 2D bounding boxes.
For objects with relatively complete 3D geometry formed by point clouds, we adopt point-cloud-clustering algorithms [6] to generate candidate 3D bounding boxes. We utilize camera parameters to project 2D bounding boxes into the point cloud space and identify the corresponding clusters. Finally, the bounding boxes are manually corrected. For objects without corresponding 3D candidates matching their 2D bounding boxes, we manually annotate 3D bounding boxes by referencing multi-view images.
For category labels, we feed images with 2D bounding boxes to an MLLM for semantic annotation and indicate for each object whether its category has been seen in each dataset. To select the best MLLM and prompts for object recognition, we manually select 30 challenging annotated image samples and evaluate the accuracy of each MLLM and prompt. We tried GPT-4V [47], Claude 3 Opus [2], and InternVL 1.5 [13], with InternVL exhibiting the best performance. Our experiments also reveal that closed image prompts, such as 2D bounding boxes or circles, yield the best results, whereas marking the object of inquiry on the image with arrows yields slightly inferior results. Objects such as open manholes and wires falling on the road are difficult to identify for existing MLLMs. This implies that OpenAD can also be utilized to test the detection and semantic recognition capabilities of MLLMs. Appendix A presents a detailed demonstration of how we improved our pipeline. The final MLLM and prompt achieve an accuracy of approximately 90% on the entire dataset.
Note that although we have utilized tools such as MLLM to automate some stages as much as possible to reduce manual workload, we have also incorporated manual verification into each stage to ensure the accuracy of each annotation.